OpenStudy (anonymous):

x - 4y = 6 3x + y = -8

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Ordered pair btw.

OpenStudy (anonymous):

Are you trying to solve a systems of equations?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@Aleah.wood1398

OpenStudy (anonymous):

okay i can help one second

OpenStudy (anonymous):

I also have two more after this.

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? -3x + 4y = -20 x + 4y = -4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (anonymous):

x=-2 y=-2 for the first question

OpenStudy (anonymous):

Are you sure? All I have is -12, 2

OpenStudy (anonymous):

what are the answers?

OpenStudy (anonymous):

(2, −1) (1, −9) (−2, −2) (−6, −3)

OpenStudy (anonymous):

o wait LOL there is 2,2

OpenStudy (anonymous):

okay let me work on the next one

OpenStudy (anonymous):

i got (4,-2) for the next one

OpenStudy (anonymous):

correcto, last one...

OpenStudy (anonymous):

Tickets to the zoo cost $14 for adults and $10 for children. A group of 12 people went to the zoo, and the tickets cost $140. The system of equations models this situation, where x is the number of adults and y is the number of children. x + y = 12 14x + 10y = 140 How many adults and how many children were in the group?

OpenStudy (anonymous):

A container holds a combination of dimes and quarters. There are 178 coins in all, and their total value is $29.80. The system of equations models this situation, where x is the number of dimes and y is the number of quarters. x + y = 178 0.1x + 0.25y = 29.8 How many dimes and quarters are in the container? This is the LAST LAST question.

OpenStudy (anonymous):

Adults=5 Children=7

OpenStudy (anonymous):

Yes thats one of the answers Your definitly getting medal'd

OpenStudy (anonymous):

Dimes=98 Quarters=80

OpenStudy (anonymous):

20%????

OpenStudy (anonymous):

Jkjk. 100% Medal'ing you rn.

OpenStudy (anonymous):

I can always try and help with anything else to!(: and hahha thanks!

OpenStudy (anonymous):

Haha np c:

OpenStudy (anonymous):

We got the same last name too :P